# Logarithmic de Rham stacks and non-abelian Hodge theory

### v1

@@ Abstract.
In this article, we introduce the logarithmic de Rham stack of a pair $(X,D),$ for $X$ a smooth variety $X$ over a field $k$ of positive characteristic $p,$ and $D$ a strict normal crossings divisor on $X.$ Using this stack, we prove a new version of logarithmic Cartier descent, and a new logarithmic non-abelian Hodge theorem for curves, both stated using a certain logarithmic Frobenius twist. Our logarithmic non-abelian Hodge theorem implies an earlier logarithmic non-abelian Hodge theorem of de Cataldo-Zhang [14].
@@ 1. Introduction
de Rham stacks in characteristic 0 were first defined by Simpson [33], as a way to ‘geometrize’ algebraic de Rham cohomology. Since then, Drinfeld [16] and Bhatt-Lurie [8] have made mixed and positive characteristic analogues of Simpson’s de Rham stack. Our definition will borrow heavily from the machinery of Bhargav-Mathew-Vologodsky-Zhang [9].
@@
Remark 1.4.
Analogously to the usual de Rham stack, we will prove in Section 5 that
where by $\operatorname{Conn}(X,D)$ we mean the category of pairs $(\mathcal{E},\nabla:\mathcal{E}\to\mathcal{E}\otimes\Omega^{1}_{X}(\log D)),$ where $\mathcal{E}\in\operatorname{Qcoh}(X)$ and $\nabla$ is a flat connection whose $p$-curvature is nilpotent.
Remark 1.5 (A sheared logarithmic de Rham stack).
In Remark 1.4, it is inconvenient for some applications to need to restrict to connections with $p$-curvature zero. To eliminate this problem, we will also define a sheared variant of $X^{\operatorname{dR}}(\log D),$ which we define by $X^{\widehat{\operatorname{dR}}}(\log D).$ The theory of sheared de Rham stacks will be reviewed in Section 2, but the key difference is that
Many connections which arise in practice have logarithmic poles; thus, for any theorem about the category $\operatorname{Conn}(X)$ (meaning flat connections on $X$ with no poles), it is often essential to have a corresponding theorem about $\operatorname{Conn}(X,D).$
The purpose of the logarithmic de Rham stack is to allow one to easily transmute the proof of a theorem about $\operatorname{Conn}(X)$ to a proof of a theorem about $\operatorname{Conn}(X,D).$ The point is that, since the logarithmic de Rham stack is defined in almost exactly the same way as the usual de Rham stack $(X/k)^{\operatorname{dR}},$ most arguments involving the de Rham stack will not care about if the base is $\operatorname{Spec}(k)$ or if it is $(\mathbb{A}^{1}/\mathbb{G}_{m})^{n}.$
@@ 1.2. Logarithmic non-abelian Hodge theory
1.2.1. Relation to work of Li-Sun
Li-Sun [25] explore a version of logarithmic non-abelian Hodge theory for curves in positive characteristic using parahoric structures, more closely analogous to the characteristic 0 story. The author is presently unaware of any direct relationship between our Theorem 1.17 and their results. While quasicoherent sheaves on root stacks (such as $(X/\Theta^{n})^{(1)}$) are often called sheaves with parabolic structure, it seems the parahoric structures used are of a very different nature than the parabolic structures considered here. Indeed, their parahoric structures depend heavily on the group used, whereas our parabolic structures are always just $p$-step filtrations.
@@ 1.3. Weight filtrations
The logarithmic de Rham stack $X^{\widehat{\operatorname{dR}}}(\log D)$ naturally lives over $(\mathbb{A}^{1}/\mathbb{G}_{m})^{n},$ since it is defined as the relative de Rham stack of a morphism over $(\mathbb{A}^{1}/\mathbb{G}_{m})^{n}.$ Thus, the cohomology of any quasicoherent sheaf on $X^{\widehat{\operatorname{dR}}}(\log D)$ is naturally equipped with a $\mathbb{Z}^{n}$-indexed filtration. We will identify this filtration with Deligne’s pole-order filtration (so, the filtered piece $F_{a_{1},...,a_{n}}$ measures the cohomology coming from differential forms which have a pole of order at most $a_{i}$ along the $i^{\text{th}}$ component of $D$). This relates to recent work of Annala-Pstragowski [2] on weight filtrations on de Rham cohomology in positive characteristic. More precisely, if $U=X\setminus D,$ then Annala-Pstragowski [2] define a $\mathbb{Z}$-indexed filtration on $H^{q}_{\operatorname{dR}}(U),$ and show it can be computed in terms of the pole-order filtration.
As Annala-Pstragowski prove that this weight filtration is independent of the choice of compactification of $U,$ it is interesting to ask the following question.
Question 1.19.
Let $U$ be a smooth variety over $k.$ Is there some stack $U^{\operatorname{dR},\operatorname{reg}}$, depending only on $U,$ such that
  1. (1)
    $\operatorname{Qcoh}(U^{\operatorname{dR},\operatorname{reg}})\simeq\operatorname{Conn}(U)^{\operatorname{reg}},$ the category of connections on $U$ with regular singularities at infinity;
  2. (2)
    and $U^{\operatorname{dR},\operatorname{reg}}$ admits a morphism $U^{\operatorname{dR},\operatorname{reg}}\to\mathbb{A}^{1}/\mathbb{G}_{m},$ such that the induced filtration on $H^{q}_{\operatorname{dR}}(U)$ is just the weight filtration of Annala-Pstragowski [2]?
@@
As a smooth curve has a unique compactification, logarithmic de Rham stacks provide a positive answer to Question 1.19 for $\dim U=1.$
@@ 1.5. Acknowledgements
The author thanks Bhargav Bhatt for many very useful discussions on this work, and for suggesting the author use these techniques to re-prove de Cataldo-Zhang’s [14] logarithmic non-abelian Hodge theorem.
In addition, the author is grateful to Niven Achenjang for explaining aspects of his work [1]; to Zachary Berens, Kenta Suzuki and Gleb Terentiuk for discussions about relative de Rham stacks over $B\mathbb{G}_{m}$; to Niels Borne for communications about [12]; to Sanath Devalapurkar for communications about $p$-curvature; to Longke Tang and Qixiang Wang for simplifying the construction of Section 3.1.2; and to Mingjia Zhang for discussions on sheared de Rham stacks.
@@ 2. The relative de Rham stack and its sheared variant
In this purely expository section, we recall a few results on de Rham stacks and sheared de Rham stacks. This is done mostly because the results we need about sheared de Rham stacks are not yet in the literature, but nothing in this section should be viewed as original; all arguments we make in this section essentially appear in Bhatt’s notes [7] (which themselves are based on the related works [17] of Drinfeld and [8] of Bhatt-Lurie). We hope this section can be useful to a reader less familiar with the theory of de Rham stacks.
@@
To reiterate, no results in this section are original; all come from Bhatt-Mathew-Vologodsky-Zhang’s forthcoming work [9], Drinfeld’s [18] and [17], and Bhatt-Lurie’s [8]. We recall these results in part because [9] has not yet appeared and in part to make it easier for the reader to know what we are citing.
@@ 2.1.1. Reminder on quasi-ideals
In an $\infty$-categorical context, the cofiber $I\to R$ is a little more natural, and the truncation can make the cone seem a little confusing to work with. We will explain that, for our examples of interest, the truncation is not truly needed.
@@
is $\tau_{\leq 1}\operatorname{cofib}(\mathcal{I}\to\mathcal{R})\simeq[\mathcal{R}/\mathcal{I}],$ as desired. ∎
@@ 2.1.3. Witt vectors and $\mathbb{G}_{a}^{\#}$
In other words, $F$ is faithfully flat and there is a $W$-linear equivalence $\mathbb{G}_{a}^{\#}\simeq W.$
@@
$\mathbb{G}_{a}^{\widehat{\operatorname{dR}}}\simeq\operatorname{cone}(F_{*}W\xrightarrow{p}F_{*}W).$
@@ 2.1.4. Sheared Witt vectors and $\mathbb{G}_{a}^{\#,\wedge}$
By Lemma 2.25, it thus suffices to prove that $R\Gamma_{\operatorname{flat}}(\operatorname{Spec}(A),W^{\prime})$ is concentrated in degree 0 to get both of our claims, since we have a short exact sequence
@@ 2.2.2. Flatness of $\pi_{X/S}$ and $\pi_{X/S,\wedge}$
Deformation theory (in derived algebraic geometry) tells us that, since $X\to\operatorname{Spec}(k)$ is smooth, and $A\to\mathbb{G}_{a}^{\widehat{\operatorname{dR}}}(A)$ is surjective on $\pi_{0}$ with nilpotent kernel, the morphism $X(A)\to X(\mathbb{G}_{a}^{\widehat{\operatorname{dR}}}(A))$ of anima is surjective on $\pi_{0}.$ As the semiperfect $k$-algebras form a basis for the flat topology, this immediately implies $\pi_{X/S,\wedge}$ is surjective in the fpqc topology. When $f$ is furthermore etale, deformation theory furthermore gives us that $X(A)\to X(\mathbb{G}_{a}^{\widehat{\operatorname{dR}}}(A))$ is an equivalence of anima; thus $X(A)\to X^{\operatorname{dR}}(A)$ is an equivalence for all semiperfect $A$, and hence $\pi_{X/S}$ is an equivalence of stacks.
@@ 2.3. The gerbe property
$^{(X^{\prime}/S^{\prime})^{\widehat{\operatorname{dR}}}}$$^{(X/S)^{\widehat{\operatorname{dR}}}}$$(X^{\prime})^{(1)}$$X^{(1)}.$$\scriptstyle{\nu_{X^{\prime}/S^{\prime},\wedge}}$$\scriptstyle{\nu_{X/S,\wedge}}$
@@
We can understand the fibers of $(X/S)^{\operatorname{dR}}\to X^{(1)}$ using deformation theory. Fix a morphism of $\mathbb{F}_{p}$-schemes $\operatorname{Spec}(A)\to S,$ and let $x:\operatorname{Spec}(A)\to X^{(1)}$ be an $A$-point of $X^{(1)}$, and let $y:\operatorname{Spec}(F_{*}A)\to X$ be the corresponding point of $X.$ An $A$-point of $(X/S)^{\operatorname{dR}}$ lifting $x$ is then the same as a map
@@
Deformation theory tells us that this fiber will either be empty, or a torsor under

### v2

@@ Abstract.
In this article, we introduce the logarithmic de Rham stack of a pair $(X,D),$ for $X$ a smooth variety over a field $k$ of positive characteristic $p,$ and $D$ a strict normal crossings divisor on $X.$ Using this stack, we prove a new version of logarithmic Cartier descent, and a new logarithmic non-abelian Hodge theorem for curves, both stated using a certain logarithmic Frobenius twist. Our logarithmic non-abelian Hodge theorem implies an earlier logarithmic non-abelian Hodge theorem of de Cataldo-Zhang [15].
@@ 1. Introduction
de Rham stacks in characteristic 0 were first defined by Simpson [39], as a way to ‘geometrize’ algebraic de Rham cohomology. That is, for $Y$ a complex algebraic variety, Simpson defined a certain stack $f:Y^{\operatorname{Simpson-dR}}\to\operatorname{Spec}(\mathbb{C})$ having the following two properties:
  1. (1)
    there is an equivalence of categories $\operatorname{Qcoh}(Y^{\operatorname{Simpson-dR}})\simeq\operatorname{Conn}(Y),$ for $\operatorname{Conn}(Y)$ the category of quasicoherent sheaves on $Y$ endowed with a flat connection,
  2. (2)
    for each $q,$ there are isomorphisms $R^{q}f_{*}\mathcal{O}_{Y^{\operatorname{Simpson-dR}}}\simeq H^{q}_{\operatorname{dR}}(Y,\mathbb{C})$ between the coherent cohomology of $Y^{\operatorname{Simpson-dR}}$ and the algebraic de Rham cohomology of $Y.$
The utility of Simpson’s construction is that now algebraic de Rham cohomology behaves as a coherent cohomology theory. Working over $Y^{\operatorname{Simpson-dR}}$ also allows one to better understand phenomena in non-abelian Hodge theory.
Since then, Drinfeld [17] and Bhatt–Lurie [9] have made mixed and positive characteristic analogues of Simpson’s de Rham stack. Recently, Bhatt–Kanaev–Mathew–Vologodsky–Zhang [8] have introduced a ‘sheared’ variant of this construction of Drinfeld and Bhatt–Lurie, which has slightly better properties. Our logarithmic de Rham stack is essentially a logarithmic variant of these constructions, which allows one to geometrize logarithmic de Rham cohomology, in the following sense.
@@
Analogously to Simpson’s de Rham stack, we will prove in Section 5 that
where by $\operatorname{Conn}(X,D)$ we mean the category of pairs $(\mathcal{E},\nabla:\mathcal{E}\to\mathcal{E}\otimes\Omega^{1}_{X}(\log D)),$ where $\mathcal{E}\in\operatorname{Qcoh}(X)$ and $\nabla$ is a flat connection whose $p$-curvature is nilpotent. It is inconvenient for some applications to need to restrict to connections with $p$-curvature zero; to eliminate this problem, we will also define a sheared variant of $X^{\operatorname{dR}}(\log D),$ which we define by $X^{\widehat{\operatorname{dR}}}(\log D).$ The theory of sheared de Rham stacks will be reviewed in Section 2, but the key difference is that
@@
Many connections which arise in practice have logarithmic poles; thus, for any theorem about the category $\operatorname{Conn}(X)$ (meaning flat connections on $X$ with no poles), it is often essential to have a corresponding theorem about $\operatorname{Conn}(X,D).$ Our logarithmic de Rham stack will allow one to easily transmute the proof of a theorem about $\operatorname{Conn}(X)$ to a proof of a theorem about $\operatorname{Conn}(X,D).$ The point is that, since the logarithmic de Rham stack is defined in almost exactly the same way as the usual de Rham stack $(X/k)^{\operatorname{dR}},$ most arguments involving the de Rham stack will not care about if the base is $\operatorname{Spec}(k)$ or if it is $(\mathbb{A}^{1}/\mathbb{G}_{m})^{n}.$
@@ 1.1. Logarithmic Cartier descent
Remark 1.14.
Lorenzon [29] has an alternative version of logarithmic Cartier descent, phrased in the language of indexed Azumaya algebras. At the present moment, the author is unaware of a direct relationship between Lorenzon’s indexed Azumaya algebras and the parabolic structures which show up in our logarithmic Cartier descent.
@@ 1.2. Logarithmic non-abelian Hodge theory
1.2.1. Relation to work of Li–Sun
Li–Sun [26] explore a version of logarithmic non-abelian Hodge theory for curves in positive characteristic using parahoric structures, more closely analogous to the characteristic 0 story. The author is presently unaware of any direct relationship between our Theorem 1.16 and their results. While quasicoherent sheaves on root stacks (such as $(X/\Theta^{n})^{(1)}$) are often called sheaves with parabolic structure, it seems the parahoric structures used are of a very different nature than the parabolic structures considered here. Indeed, their parahoric structures depend heavily on the group used, whereas our parabolic structures are always just $p$-step filtrations.
1.2.2. Relation to work of Ogus–Vologodsky and Schepler
Ogus–Vologodsky [34] prove a different sort of logarthmic non-abelian Hodge theorem; Schepler [38] found a logarithmic variant of their work. The key differences between our work and Schepler is that
  1. (1)
    we give a logarithmic version of the non-abelian Hodge theorem of Groechenig [21], which gives a correspondence between all Higgs fields and all connections, but only in dimension 1;
  2. (2)
    Schepler [38] gives a logarithmic version of Ogus–Vologodsky’s [34] non-abelian Hodge theorem, which gives a correspondence in all dimensions, but only for Higgs fields and connections obeying a certain nilpotence condition.
Gleb Terentiuk is currently writing an article on Ogus–Vologodsky’s work, but re-done in the language of de Rham stacks; it seems very likely to the author of this work that Terentiuk’s results will also give a strengthening Schepler’s [38] result.
@@ 1.3. Weight filtrations
The logarithmic de Rham stack $X^{\widehat{\operatorname{dR}}}(\log D)$ naturally lives over $(\mathbb{A}^{1}/\mathbb{G}_{m})^{n},$ since it is defined as the relative de Rham stack of a morphism over $(\mathbb{A}^{1}/\mathbb{G}_{m})^{n}.$ Thus, the cohomology of any quasicoherent sheaf on $X^{\widehat{\operatorname{dR}}}(\log D)$ is naturally equipped with a $\mathbb{Z}^{n}$-indexed filtration. In a forthcoming work, the author will focus more on these aspects of the logarithmic de Rham stack (which are interesting even in characteristic 0).
For now, we pose the following questions. Recently, Annala-Pstragowski [2] constructed filtrations on logarithmic de Rham cohomology in positive characteristic; they then proved that these filtrations actually depend only on the open part $U:=X\setminus D,$ and not on the specific choice of compactification.
Question 1.18.
Can one geometrize the weight filtration of Annala-Pstragowski?
@@
The logarithmic de Rham stack $X^{\operatorname{dR}}(\log D)$ does not depend on just $X\setminus D,$ so in searching for a geometrization of this weight filtration of Annala-Pstragowski, it would be nice to find some modification of the logarithmic de Rham stack which does only depend on $X\setminus D.$
In this vein, it is also interesting to ask the following question.
Question 1.20.
Let $U$ be a smooth variety over $k.$ Is there some stack $U^{\operatorname{dR},\operatorname{reg}}$, depending only on $U,$ such that $\operatorname{Qcoh}(U^{\operatorname{dR},\operatorname{reg}})\simeq\operatorname{Conn}(U)^{\operatorname{reg}},$ the category of connections on $U$ with regular singularities at infinity?
Remark 1.21.
As a smooth curve has a unique compactification, logarithmic de Rham stacks provide a positive answer to Question 1.20 for $\dim U=1.$
@@ 1.5. Acknowledgements
The author thanks Bhargav Bhatt for many very useful discussions on this work, and for suggesting the author use these techniques to re-prove de Cataldo–Zhang’s [15] logarithmic non-abelian Hodge theorem.
The author also thanks Siqing Zhang for giving very detailed feedback on a preliminary draft of this paper, and for catching some typos in the first draft.
In addition, the author is grateful to Toni Annala for explaining aspects of [2]; to Niven Achenjang for explaining aspects of his work [1]; to Zachary Berens, Kenta Suzuki and Gleb Terentiuk for discussions about relative de Rham stacks over $B\mathbb{G}_{m}$; to Niels Borne for communications about [13]; to Sanath Devalapurkar for communications about $p$-curvature; to Longke Tang and Qixiang Wang for simplifying the construction of Section 3.1.2; to Siqing Zhang for giving detailed feedback on a draft of this article; and to Mingjia Zhang for discussions on sheared de Rham stacks.
@@ 2. The relative de Rham stack and its sheared variant
In this purely expository section, we recall a few results on de Rham stacks and sheared de Rham stacks. This is done mostly because the results we need about sheared de Rham stacks are not yet in the literature, but nothing in this section should be viewed as original; all arguments we make in this section essentially appear in Bhatt’s notes [7] (which themselves are based on the related works [18] of Drinfeld and [9] of Bhatt–Lurie). We hope this section can be useful to a reader less familiar with the theory of de Rham stacks.
@@
To reiterate, no results in this section are original; all come from Bhatt–Mathew–Vologodsky–Zhang’s forthcoming work [10], Drinfeld’s [19] and [18], and Bhatt–Lurie’s [9]. We recall these results in part because [10] has not yet appeared publicly and in part to make the article a little more self-contained.
@@ 2.1.1. Reminder on quasi-ideals
In an $\infty$-categorical context, the cofiber $\operatorname{cofib}(I\to R)$ is a little more natural, and the truncation can make the cone seem a little confusing to work with. We will explain that, for our examples of interest, the truncation is not truly needed.
@@
is $\tau_{\leq 1}\operatorname{cofib}(\mathcal{I}\to\mathcal{R})\simeq[\mathcal{R}/\mathcal{I}],$ as desired. This last identification with the quotient stack is because the cofiber, computed in the category of presheaves, has functor of points
$\operatorname{cofib}(\mathcal{I}(T)\to\mathcal{R}(T)),$
whose 1-truncation is just the groupoid whose objects are elements of the ring $\mathcal{R}(T),$ and where morphisms $r\to r^{\prime}$ are just the data of some $i\in\mathcal{I}(T)$ so that $r^{\prime}=r+i.$ Sheafifying this presheaf of groupoids produces the usual quotient stack. ∎
@@ 2.1.3. Witt vectors and $\mathbb{G}_{a}^{\#}$
In other words, $F$ is faithfully flat and there is a $W$-linear equivalence $\mathbb{G}_{a}^{\#}\simeq W[F].$
@@
$\mathbb{G}_{a}^{\operatorname{dR}}\simeq\operatorname{cone}(F_{*}W\xrightarrow{p}F_{*}W).$
@@ 2.1.4. Sheared Witt vectors and $\mathbb{G}_{a}^{\#,\wedge}$
By Lemma 2.25, it thus suffices to prove that $R\Gamma_{\operatorname{flat}}(\operatorname{Spec}(A),W^{\prime})$ is concentrated in degree 0 to get our claim, since we have a short exact sequence
@@ 2.2.2. Flatness of $\pi_{X/S}$ and $\pi_{X/S,\wedge}$
Before giving our result, we recall a basic result from derived deformation theory.
Lemma 2.37.
Let $X\to\operatorname{Spec}(k)$ be a $k$-scheme, and $A$ an animated $k$-algebra. If $B$ is a square-zero extension of $A$ by $M,$ then for any point $\eta\in X(A),$ there is a fiber sequence of anima
$\{\eta\}\times_{X(A)}X(B)\to\operatorname{pt}\to\operatorname{Map}_{A}(\eta^{*}\mathbb{L}_{X/k},M[1]).$
Remark 2.38.
The middle term is $\operatorname{pt}$ because we’ve fixed $k$-algebra structures on $A,B.$
Proof.
Lurie [32] proves the spectral algebraic geometry version of this in his Remark 17.3.9.2, but the proof goes through for animated rings; for the reader’s benefit, we reproduce it.
By definition of square-zero extension of animated rings, there is a pullback square of animated rings
[…]
where $i$ is the inclusion and $s$ is a certain derivation.
By Lurie’s [30] Proposition 5.3.7, the functor of points of a scheme is infinitesimally cohesive, which implies that the induced diagram
[…]
is a Cartesian diagram in $\operatorname{Ani}.$
The projection $A\oplus M[1]\to A$ induces a map
$X(A\oplus M[1])\to X(A).$
By definition of the cotangent complex, we have
$\operatorname{Map}_{A}(\eta^{*}\mathbb{L}_{X/k},M[1])\simeq X(A\oplus M[1])\times_{X(A)}\{\eta\}.$
Taking the fiber above $\eta$ in our $X(B)$ Cartesian diagram, we then get a Cartesian diagram
$\{\eta\}\times_{X(A)}X(B)$$\operatorname{pt}$$\operatorname{pt}$$\operatorname{Map}_{A}(\eta^{*}\mathbb{L}_{X/k},M[1]),$}
as desired. ∎
Corollary 2.39.
Suppose $X\to\operatorname{Spec}(k)$ is a smooth $k$-scheme, and $B\to A$ is a morphism of animated $k$-algebras such that $\pi_{0}(A)\to\pi_{0}(B)$ is surjective with nilpotent kernel. Then $X(B)\to X(A)$ is surjective on $\pi_{0}.$ If $X$ is moreover etale, then $X(B)\to X(A)$ is an equivalence.
Proof.
As in Remark 3.4.2 of Lurie’s [30], it suffices to treat the case of square-zero extensions, so assume $B$ is a square-zero extension of $A$ by some $M.$
For any $\eta\in X(A),$ we have by Lemma 2.37 a fiber sequence
$\{\eta\}\times_{X(A)}X(B)\to\operatorname{pt}\to\operatorname{Map}_{A}(\eta^{*}\mathbb{L}_{X/k},M[1]).$
When $X$ is smooth, $\mathbb{L}_{X/k}$ lives in degree 0, so that
$\pi_{0}(\operatorname{Map}_{A}(\eta^{*}\mathbb{L}_{X/k},M[1]))=\operatorname{pt},$
and hence $\pi_{0}(\{\eta\}\times_{X(A)}X(B))=\operatorname{pt},$ implying the claimed surjectivity.
When $X$ is moreover etale, we have $\mathbb{L}_{X/k}=0,$ so that in fact every fiber of $X(A)\to X(B)$ has the homotopy type of a point. It then follows from Whitehead’s theorem that $X(A)\to X(B)$ is an equivalence in this situation, as desired. ∎
@@
Deformation theory (see Corollary 2.39) tells us that, since $X\to\operatorname{Spec}(k)$ is smooth, and $A\to\mathbb{G}_{a}^{\widehat{\operatorname{dR}}}(A)$ is surjective on $\pi_{0}$ with nilpotent kernel, the morphism $X(A)\to X(\mathbb{G}_{a}^{\widehat{\operatorname{dR}}}(A))$ of anima is surjective on $\pi_{0}.$ As the semiperfect $k$-algebras form a basis for the flat topology, this immediately implies $\pi_{X/S,\wedge}$ is surjective in the fpqc topology. When $f$ is furthermore etale, deformation theory (again see Corollary 2.39) furthermore gives us that $X(A)\to X(\mathbb{G}_{a}^{\widehat{\operatorname{dR}}}(A))$ is an equivalence of anima; thus $X(A)\to X^{\operatorname{dR}}(A)$ is an equivalence for all semiperfect $A$, and hence $\pi_{X/S}$ is an equivalence of stacks.
@@ 2.3. The gerbe property
$^{(X^{\prime}/S^{\prime})^{\widehat{\operatorname{dR}}}}$$^{(X/S)^{\widehat{\operatorname{dR}}}}$$(X^{\prime})^{(1)}$$X^{(1)}$$\scriptstyle{\nu_{X^{\prime}/S^{\prime},\wedge}}$$\scriptstyle{\nu_{X/S,\wedge}}$
commutes.
@@
We can understand the fibers of $(X/S)^{\operatorname{dR}}\to X^{(1)}$ using deformation theory (more particularly, in the form of Lemma 2.37). Fix a morphism of $\mathbb{F}_{p}$-schemes $\operatorname{Spec}(A)\to S,$ and let $x:\operatorname{Spec}(A)\to X^{(1)}$ be an $A$-point of $X^{(1)}$, and let $y:\operatorname{Spec}(F_{*}A)\to X$ be the corresponding point of $X.$ An $A$-point of $(X/S)^{\operatorname{dR}}$ lifting $x$ is then the same as a map
@@
Deformation theory (see Lemma 2.37) tells us that this fiber will either be empty, or a torsor under